To prevent the presence of air, noble gases are placed over highly reactive chemicals to act as inert blanketing gases. A chemical engineer places a mixture of noble gases consisting of 5.50g of He, 15.0g of Ne and 35.0g Kr in a piston cylinder assembly at STP=0 degree Celsius and 1atm. RAM: He 4.003g/mole, Ne 20.18g/mole and Kr 83.80g/mole. What is the partial pressure for each noble gas?
1
Expert's answer
2018-11-27T06:59:41-0500

Answer on Question #83331, Chemistry/ General Chemistry

To prevent the presence of air, noble gases are placed over highly reactive chemicals to act as inert blanketing gases. A chemical engineer places a mixture of noble gases consisting of 5.50g of He, 15.0g of Ne and 35.0g Kr in a piston cylinder assembly at STP=0 degree Celsius and 1atm. RAM: He 4.003g/mole, Ne 20.18g/mole and Kr 83.80g/mole. What is the partial pressure for each noble gas?

Solution

Dalton's Law of partial pressure

Ptotal = P(He) + P(Ne) + P(Kr)

P(He) = X(He)*Ptotal, P(Ne) = X(Ne)*Ptotal, P(Kr) = X(Kr)*Ptotal

X(He)= n(He)/ntotal, X(Ne)= n(Ne)/ntotal , X(Kr)= n(Kr)/ntotal

ntotal = n(He) + n(Ne) + n(Kr)

Calculations:

n= m/M

n(He) = 5.50 g/ 4.003 g/mole = 1.374 mol

n(Ne) = 15.0 g / 20.18g/mole = 0.7433 mol

n(Kr) = 35.0 g/83.80 g/mole = 0.4177 mol

ntotal = 1.374 mol + 0.7433 mol + 0.4177 mol = 2.535 mol

P(He) = X(He)*Ptotal = (1.374 mol/2.535mol)(1 atm) = 0.5420 atm

P(Ne) = X(Ne)*Ptotal = (0.7433 mol/ 2.535 mol)(1 atm) = 0.2932 atm

P(Kr) = X(Kr)*Ptotal = (0.4177 mol/2.535 mol)*(1 atm) = 0.1648 atm

Answer: P(He) = 0.5420 atm

P(Ne) = 0.2932 atm

P(Kr) = 0.1648 atm